a) Calculate the amount of chlorine gas evolved during the electrolysis of molten potassium chloride on passing a current of 0.5 A for two hours through it.


b) Define a potentiometric titration and draw a schematic graph of a redox potentiometric titration.


1
Expert's answer
2021-03-17T06:44:10-0400

( a) 2Cl- = Cl2 + 2e


For 2 Faraday's 1 mol of Cl2 evolved .

So , 0.5 Faraday's 0.25 mol of Cl2 gas evolved.


Mass of Cl2 gas = 0.25 × 71 = 17.75 g .


( b ) Potentiometric titration is a technique similar to direct titration of a redox reaction. It is a useful means of characterizing an acid. No indicator is used; instead the potential is measured across the analyte, typically an electrolyte solution.